### v2

# HandSSCA: 3D Hand Mesh Reconstruction with State Space Channel Attention from RGB images

Abstract
Reconstructing a hand mesh from a single RGB image is a challenging task because hands are often occluded by objects. Most previous works attempted to introduce more additional information and adopt attention mechanisms to improve 3D reconstruction results, but it would increased computational complexity. This observation prompts us to propose a new and concise architecture while improving computational efficiency. In this work, we propose a simple and effective 3D hand mesh reconstruction network HandSSCA, which is the first to incorporate state space modeling into the field of hand pose estimation. In the network, we have designed a novel state space channel attention module that extends the effective sensory field, extracts hand features in the spatial dimension, and enhances hand regional features in the channel dimension. This design helps to reconstruct a complete and detailed hand mesh. Extensive experiments conducted on well-known datasets featuring challenging hand-object occlusions (such as FREIHAND, DEXYCB, and HO3D) demonstrate that our proposed HandSSCA achieves state-of-the-art performance while maintaining a minimal parameter count.
1 Introduction
Reconstruction of 3D hand pose from a single RGB image has a wide range of applications in many fields, such as VR/AR, robotics, and human-computer interaction. In order to reduce the cost of use in above areas, 3D hand pose estimation based on RGB images has attracted much attention due to its low cost and computational friendliness, which provides new solutions for practical applications. Influenced by the rapid development of deep learning, many excellent works Hasson2019LearningJR; Boukhayma20193DHS; Choutas2020MonocularEB; Hampali2021KeypointTS; Wang_2023; park2022handoccnet; li2022interacting have emerged in this field. These works focus on issues related to hand-object occlusion and the loss of local information. Despite significant advancements, most methods still struggle to reconstruct hand poses in scenarios with substantial occlusion or when there is a lack of prior knowledge.
The previous methods can be broadly categorized into two groups, one is by incorporating additional a priori knowledge. For example, Zimm et al. zimmermann2017learning obtained hand region features via a hand segmentation network. And Wang et al. Wang_2023 added keypoint coordinates as strong a priori information to constrain hand pose. These methods can enhance pose estimation performance, but they lack essential a priori information in practical applications. Another category is to learn the intrinsic connection between image features through deep networks, Park et al. park2022handoccnet extracts the complete hand information by distinguishing the features. jiao et al. https://doi.org/10.1049/ell2.13040 uses the Transformer architecture to obtain the global information and expects to learn the hand feature mapping from the occluded part. These methods tend to mine the association between image features, but there are still some limitations. First, the Attention mechanism leads to excessive computational complexity. Second, an overload of global information can interfere with the effective receptive field luo2017understanding, as shown in Figure.1(b).
Based on the above observations, we propose a simple and effective 3D hand reconstruction network HandSSCA. This framework for the first time incorporates state space model liu2024vmamba; ma2024umamba; ruan2024vmune into the field of hand pose estimation. At the same time, combining the idea of spatial channel feature fusion hu2019squeezeandexcitation, State Space Channel Attention (SSCA) was constructed, which emphasizes the contribution of channel dimension features to image features, as shown in Fig.3. This not only increases the range of effective receptive fields in the network, but also effectively avoids the computational complexity of traditional attention square steps.Specifically, our main contributions are as follows:
(1) We propose a hand pose estimation model called HandSSCA, which introduces the concept of state space model to the field of hand pose estimation for the first time. This network can effectively improve hand reconstruction performance without the need for additional prior knowledge.
(2) We propose a spatial and channel-based parallel scanning approach to compensate for the missing channel information in planar scanning. And the state space channel attention module is constructed, which can enhance the effective receptive field range while maintaining a small number of parameters.
(3) Our method achieves state-of-the-art on three datasets, FREIHAND, DEXYCB and HO3D, with a small number of parameters.
Figure 1: Effective receptive fields (from left to right, FPN, Self-Attention, State Space Attention(SSA) and State Space Channel Attention(SSCA), respectively).
2 Related work
2.1 State Space Model
With the proposal of attention mechanism vaswani2023attention, it makes more models can effectively focus on global information mapping, but its quadratic computational complexity leads to computational inefficiency in the case of long sequences. In the field of natural language processing, Albert Gu et al. gu2022efficiently proposed the S4 model to optimize the state space model (SSM) by simulation in order to solve the problems of computational efficiency and long sequence dependency. However, the S4 model would process all inputs in exactly the same way, which is not in line with the practical intuition that each sequence should receive different attention. Therefore, Albert Gu et al. gu2023mamba introduced a selection mechanism to propose the Mamba architecture (S6) to learn the parameters in the traditional state-space model, through which it is able to focus on or ignore some sequences. Compared to traditional attention computation methods, S6 enables each element in a 1-D array (e.g., a text sequence) to interact with any previously scanned samples through a compressed hidden state, effectively reducing the quadratic complexity to linear. Influenced by this idea, Yue Liu et al. ma2024umamba introduced it to vision by combining the VIT model dosovitskiy2021image to transform images into sequences, and proposed a cross-scanning module to solve the problem of image orientation sensitivity. This approach achieves linear complexity without sacrificing the global receptive field, as shown in Figure.1(c). In addition, some researchers ruan2024vmunet; hu2019squeezeandexcitation have also combined state-space modeling with traditional CNN architectures to extract multi-scale global features in U-shaped structures by state-space selection modeling, and retained the deep-level information.
2.2 Additional information based Methods
It is difficult to restore 3D hand gestures from RGB image features alone, and most researchers use heatmaps Boukhayma20193DHS; lv2021handtailor; iqbal2018hand, segmentation zimmermann2017learning; zhang2021hand, depth maps rezaei2022trihornnet; rezaei2023pushing; 10.1007/978-3-030-01270-0_15, and density maps zhao2023semisupervised; chen2021i2uvhandnet to incorporate additional information to help the network better learn hand features in occlusion situations. In recent years, some scholars Wang_2023; jiang2023a2jtransformer combined human a priori knowledge to guide the network to focus on the target region, Wang et al. Wang_2023 proposed HandGCAT, which learns the a priori knowledge of the hand joint positions through the Knowledge-Guided Graph Convolution module, and then through the Cross-Attention Transformer to guide the network to inject joint information into the occluded region. In addition, other scholars Hasson2019LearningJR; wang2022interacting; qi2024hoisdf; li2022artiboost; chen2022tracking have argued that the object pose affects the hand pose, therefore, Qi et al. qi2024hoisdf introduced Signed Distance Field (SDF) to reconstruct the occluded object and the hand, and this method is able to determine the grip state of the hand by estimating the object pose and then determining the hand’s grasping state.
2.3 Attention based Methods
Attentional mechanisms have achieved great success in the field of deep learning, thanks to their ability to map long-term contextual relevance, and in the task of hand pose estimation, occluded hand features can be learned through global mapping, which in turn can effectively solve the occlusion problem park2022handoccnet; li2022interacting; cho2022crossattention; feng2021collaborative; pavlakos2023reconstructing. In establishing the global context of hand-object interaction, Pavlakos et al. pavlakos2023reconstructing learn the mapping relationship between hand and occlusion from large-scale data through the VIT architecture.Cho et al. cho2022crossattention obtain the features related to the region of interest from the occluded portion through cross-attention. Park et al. park2022handoccnet make full use of the information of occluded region as an enhancement of image feature of the hand and refine it to get the complete hand features.
Figure 2: Overall architecture of HandSSCA including Trunk, SSCA and Regressor.Our HandSSCA uses Resnet50 to extract image features and fuses multiscale features by upsampling layer by layer. It then uses SSCA to obtain hand region features by pairwise spatial and channel scanning. Finally, the regressor obtains keypoint information by heatmap encoding and outputs MANO parameters. The final 3D hand mesh is obtained by forwarding the MANO parameters to the MANO layer.
3 Method
3.1 Overview
The overall network architecture is shown in Figure.2 and consists of Backbone, SSCA and Regressor components. Given a hand image $I\in\mathbb{R}^{256\times 256\times 3}$. First, the image information is extracted by ResNet-50 and the multi-scale features are fused layer by layer to get the feature mapping map $F\in\mathbb{R}^{256\times 32\times 32}$. Then, it is fed into the SSCA module and scanned in both spatial and channel dimensions to compute a 2D feature map and fused to obtain hand features $H\in\mathbb{R}^{256\times 32\times 32}$. Finally, the hand features $H$ are passed to Regressor to calculate the keypoint heat map $HM\in\mathbb{R}^{21\times 32\times 32}$ to predict the keypoint coordinates and output the MANO parameters through the FC layer to obtain the final 3D hand mesh.
3.2 Backbone
In the feature extraction part we use ResNet-50 he2015deep to extract image information and fuse multi-scale features layer by layer lin2017feature. Specifically, when the input image size is $256\times 256\times 3$, the size of the first layer of features after ResNet-50 becomes $128\times 128\times 128$. In the subsequent layers, each layer of image features $F_{i}\in\frac{H_{i}}{2^{i}}\times\frac{W_{i}}{2^{i}}\times C_{i},i\in(1,2,3,4),C_{i}\in(256,512,1024,2048)$ is obtained by downsampling. Then, up-sampling is performed and the image features $F_{i}$ are incorporated layer by layer to get the last layer of features of size $64\times 64\times 256$. Finally, the feature mapping map $F\in\mathbb{R}^{256\times 32\times 32}$ is obtained by fusing the image information through average pooling.
3.3 State Space Channel Attention
State Space Models.
S4ND nguyen2022s4nd was the first study to introduce state space models into the field of vision by enabling it to receive 2D image information through a simple extension of the S4 model. The S6 module proposed by Gu et al. gu2023mamba was shown in VMamba liu2024vmamba to be able to capture the global feature information of 2D image features through a selectable scanning mechanism enabling it to capture the global feature information of the 2D image features and maintain linear complexity. The state space model expects to map the input $x(t)\in\mathbb{R}^{L}$ to the output $y(t)\in\mathbb{R}^{L}$ through the implicit state $h(t)$ and the parameter $A\in\mathbb{C}^{N\times N},B,C\in\mathbb{C}^{N},D\in\mathbb{C}^{1}$ with the following equation:
$h^{{}^{\prime}}(t)=Ah(t)+Bx(t),\\ \mathrm{y(t)=Ch(t)+Dx(t)}$ (1)
In order to achieve computationally efficient operations zimmermann2019freihand, the above ordinary differential equations need to be transformed into discrete functions. Specifically, when the input size is $X_{i}\in\mathbb{R}^{\mathrm{L\times D}}$, the discretization is performed following the following rule:
$h_{i}=\overline{A}h_{i-1}+\overline{B}x_{i},\\ y_{i}=\overline{C}h_{i}+\overline{D}x_{i},\\ \overline{A}=e^{\Delta A},\\ \overline{B}=(e^{\Delta A}-I)A^{-1}B,\\ \overline{C}=C\$ (2)
where $\mathrm{B,C\in\mathbb{C}^{D\times N}}$ and $\Delta\in\mathbb{R}^{\mathrm{D}}$. In practice, we improve the $\bar{B}$ approximation by means of a first-order Taylor series ref16:
$B=(e^{\Delta A}-I)A^{-1}B\approx(\Delta A)(\Delta A)^{-1}\Delta B=\Delta B$ (3)
Unlike traditional spatial state models, the S6 block introduces a selective scanning mechanism, where $B\in\mathbb{R}^{B\times L\times N}\quad,C\in\mathbb{R}^{B\times L\times N}$, and $\Delta\in\mathbb{R}^{B\times L\times D}$ can be obtained from inputs of size $X\in\mathbb{R}^{B\times L\times D}$. this means that S6 can sense contextual information from the inputs and update the weights.
Space select.
Unlike textual information, image information contains 2D spatial information such as local texture and global structure, and it is difficult to obtain the contextual relationship of the local sense field of an image by scanning with a simple S6 module. The 1-D convolution is extended to 2-D convolution in S4ND gu2021combining, and this approach leads to the loss of the ability to dynamically update weights based on context. In VMamba liu2024vmamba, rows and columns are unfolded and scanned along in four different directions: top-left to bottom-right, bottom-right to top-left, top-right to bottom-left, and bottom-left to top-right, by means of a cross-scanning module. This allows any pixel to contain contextual information from all four directions and can enhance the size of the sensory field in the $H\times W$ plane. Then, the sequences from each of the four directions are passed through the S6 block to compute the feature map. Finally, the four sequences are merged to obtain a new sequence of the same size as the input.
Channel select.
Although spatial features such as local texture and global structure intrinsic to 2D image features are already available through Liu et al liu2024vmamba. However, in deep image features, channels contain richer and diverse contextual information, which is mostly ignored by these methods. Therefore, we design a spatial and channel-based parallel scanning approach to compensate for the missing channel information in planar scanning, as shown in Figure.3. We perform scan expansion and scan fusion on the $H\times C$ plane. First, the input features are expanded on the $H\times C$ plane and scanned from four directions to obtain the sequence $C_{1},C_{2},C_{3},C_{4}$. Then, they are fed into the S6 block to compute the contextual correlation in the sequence by Eq. (3), respectively. Finally, the sequences are merged and reconstructed into image feature maps of the same size.
Image features are spatially selected by activating the mapping of each pixel in the spatial plane and the four directions of up, down, left and right to obtain local texture and structure information. In channel selection, the mapping relationship between different channels can be obtained to enhance the representation of similar channels in the implicit space in order to construct a wide range of global feature mappings and to enhance the scope of the sensory field. We visualized the range of effective receptive fields for different attentions, as shown in Figure.1. As can be seen in Figure.1(d), this novel state space channel attention is able to maintain a larger and deeper receptive field range and global feature mapping capability with a reduced number of parameters and linear complexity. This is consistent with our motivation to expect that enhancing the effective receptive field can better capture hand-object information associations from image features.
Figure 3: SSCA module framework, we send the image features to parallel spatial and channel branches for scanning, and then process the spatial and channel resultant features and merge them through S6 block respectively, and finally sum them to get the 2D feature map.
3.4 Regressor
In Regressor, we input the hand attention features obtained from SSMA. First, a 2D heat map of each joint $H$ is output through a simple Conv cascade. Then, a cascade of hand attention features and 2D heatmap $HM$ is input into four residual blocks [34]. Finally, the outputs of the residual blocks are vectorized into 2048-dimensional vectors and passed to the fully-connected layer for predicting the pose parameter $\theta\in\mathbb{R}^{48}$ and shape parameter $\mathfrak{\beta}\in\mathbb{R}^{10}$ of the MANO. At the same time, the joint regression matrices are multiplied by the original 3D mesh $\overline{\bar{T}}$ , and the forward kinematics are applied to obtain the final 3D hand joint coordinates $J\in\mathbb{R}^{21\times 3}$ and the 3D hand mesh $V\in\mathbb{R}^{778\times 3}$.
To train our proposed HandSSCA hand pose estimation model, we constrain the model by calculating L2 loss on MANO parameters, 2D hand keypoints, 3D hand keypoints and vertex coordinates.
$\mathrm{Loss}=\alpha_{\theta}\|\theta^{\prime}-\theta\|+\alpha_{\beta}\|\beta^{\prime}-\beta\|+\alpha_{2d}\|J_{2d}^{\prime}-J_{2d}\|+\alpha_{3d}\|J_{3d}^{\prime}-J_{3d}\|+\alpha_{V}\|V^{\prime}-V\|$ (4)
Where $*^{\prime}$ denotes the true value, * denotes the predicted value, and $\alpha_{\theta}=10,\alpha_{\beta}=0.1,\alpha_{2d}=1\mathrm{e}2,\alpha_{3d}=1\mathrm{e}4,\alpha_{V}=1\mathrm{e}4$. the optimal solution is obtained by minimizing this objective function for training.
Table 1: COMPARISON WITH STATE-OF-THE-ART METHODS ON FREIHAND.
Method Params↓ PA-MPJPE↓ PA-MPVPE↓ F@5↑ F@15↑
Hasson et al.Hasson2019LearningJR 33.5M 13.3 13.3 0.429 0.907
ExPoseChoutas2020MonocularEB 11.8 12.2 0.484 0.918
Kulon et al.kulon2020weaklysupervised 519.0M 8.4 8.6 0.614 0.966
I2L-MeshNetmoon2020i2lmeshnet 136.8M 7.4 7.6 0.681 0.973
Pose2Meshchoi2021pose2mesh 75.0M 7.4 7.6 0.683 0.973
Tang et al.9710830 156.5M 6.7 6.7 0.724 0.981
Junhyeong et al.cho2022crossattention 153.0M 6.5 0.982
HandSSCA(Ours) 31.8M 7.6 7.0 0.690 0.982
Table 2: COMPARISON WITH STATE-OF-THE-ART METHODS ON DEXYCB.
Method Params↓ MPJPE↓ PA-MPJPE↓
Spurr et al.spurr2020weakly 17.34 6.83
METROlin2021endtoend 229.5M 15.24 6.99
Liu et al.liu2021semisupervised 34.9M 15.28 6.58
HandOccNetpark2022handoccnet 39.3M 14.04 5.80
HandGCATWang_2023 75.8M 13.76 5.60
HandSSCA(Ours) 31.8M 12.81 5.52
Table 3: COMPARISON WITH STATE-OF-THE-ART METHODS ON HO3D.
Method Params↓ PA-MPJPE↓ PA-MPJPE AUC↑ PA-MPVPE↓ PA-MPVPE AUC↑ F@5↑ F@15↑
I2L-MeshNetmoon2020i2lmeshnet 136.8M 11.2 0.775 13.9 0.722 0.409 0.932
Hasson et al.hasson2020leveraging 13.2M 11.0 0.780 11.2 0.777 0.464 0.939
Hampali et al.hampali2020honnotate 10.7 0.788 10.6 0.790 0.506 0.942
METROlin2021endtoend 229.5M 10.4 0.792 11.1 0.779 0.484 0.946
Liu et al.liu2021semisupervised 34.9M 9.9 0.803 9.5 0.810 0.528 0.956
I2UV-HandNetchen2021i2uvhandnet 9.9 0.804 10.1 0.799 0.500 0.943
ArtiBoostli2022artiboost 25.3M 11.4 0.773 10.9 0.782 0.488 0.944
KeypointHampali2021KeypointTS 56.5M 10.8 0.786
HandSSCA(Ours) 31.8M 9.5 0.808 9.5 0.810 0.518 0.955
4 Experiments
4.1 Implementation details
All implementations were done in PyTorch. We used the Adam optimizer kingma2017adam, and each model was trained by annealing every 10th time starting from an initial learning rate of 10-4 learning rate to train with a batch size of 32. Both our training and testing were performed on a server with a CPU15 vCPU AMD and an RTX A5000 (24GB). All other details will be provided in our code.
4.2 Datasets and evaluation metrics
FREIHAND.
This dataset zimmermann2019freihand presents the first massively diverse hand dataset containing 32,560 training samples with the possibility of replacing four backgrounds. Also, there are 3960 test samples under occlusion and poor lighting.
DEXYCB.
This dataset chao2021dexycb is a large-scale dataset for object grasping and contains 582K image frames grasped on 20 YCB objects. We use the official “S0” segmentation to evaluate the right hand pose.
HO3D.
This dataset hampali2020honnotate is a hand-object interaction dataset with severe occlusion and contains 66,034 training samples and 11,524 test samples. The dataset provides RGB images with MANO-based hand joints and meshes and camera parameters. The test set provides only the annotations of the RGB images and the hand bounding box, and the test results can be evaluated through an online submission system.
Evaluation metrics.
For FreiHand, we report PA-MPJPE and PA-MPVPE in mm as well as F-scores. For HO3D we report PA-MPJPE/MPVPE, AUC and F-scores returned from the official evaluation server. For DexYCB we report MPJPE and PA-MPJPE in mm.
Table 4: ABLATION RESULTS FOR THE FREIHAND.
Method PA-MPJPE↓ PA-MPVPE↓ F@5↑ F@15↑
Baseline 8.0 7.4 0.669 0.977
Baseline+Self-Attention 7.7 7.0 0.688 0.981
Baseline+SSCA 7.6 7.7 0.690 0.982
Table 5: ABLATION RESULTS FOR THE DEXYCB.
Method MPJPE↓ PA-MPJPE↓
Baseline 14.07 5.78
Baseline+Self-Attention 14.40 5.81
Baseline+SSCA 12.81 5.52
Table 6: ABLATION RESULTS FOR THE HO3D.
Method PA-MPJPE↓ PA-MPVPE↓ F@5↑ F@15↑
Baseline 9.8 9.7 0.516 0.954
Baseline+Self-Attention(4.6M) 10.1 10.1 0.502 0.947
Baseline+SSA(0.57M) 9.8 9.8 0.521 0.952
Baseline+SSCA(0.58M) 9.5 9.5 0.518 0.955
4.3 Comparisons with the State-of-the-art Methods
We compare our proposed method with several recent state-of-the-art methods on three datasets, FreiHand, DEXYCB, and HO3D, as shown in Tables 1-3. As can be seen from Table 1, our method maintains state-of-the-art performance in the F@15 metric on a smaller number of datasets, although it is slightly worse than the other methods in the PA-MPJPE metric cho2022crossattention; 9710830. Moreover, its number of parameters shrinks by a factor of nearly 5 compared to models with higher accuracy cho2022crossattention. On the HO3D dataset, our method reduces the number of parameters by about 40% compared to recent methods Hampali2021KeypointTS. Meanwhile, when the number of parameters is similar to the method of Liu et al. liu2021semisupervised, all other metrics are greatly improved, and only the F-score metric is slightly reduced. To verify the effectiveness of our proposed method in the case of large-scale occlusion, the metrics on the DEXYCB dataset are shown in Table 2, and all the metrics outperform the recent state-of-the-art models Wang_2023. In conclusion, the experimental results on all three datasets show that our proposed method can still be effective in estimating hand postures when a large number of occlusions are included. Moreover, the state-of-the-art performance can be maintained while reducing the number of parameters, as shown in Figure.4.
4.4 Ablation Study
In order to validate the effectiveness of the HandSSCA method, we further analyzed the different attention modules. The results of the ablation experiments on different datasets are shown in Tables 4-6. On the FreiHand and DEXYCB datasets, we designed three experiments, Baseline, Adding Self-Attention Module, and Adding SSCA Module, as a way to validate that our proposed SSCA is able to expand the effective receptive field while enhancing the hand features. On the HO3D dataset, we designed four experiments, which are Baseline, adding the self-attention module, adding the traditional state space attention module (SSA) and adding the state space channel attention module (SSCA). In this way, we demonstrate that the state space model with the addition of channel information can enhance the ability to extract hand features in the occluded state while having a wider effective receptive field.
From Tables 4 and 5, we can see that in the presence of occlusion on the hand, the global feature information obtained from the self-attention mechanism, although it can enhance the hand features to a certain extent, will introduce more noise information. In Figure.5, we visualize the attention distribution map, and it can be clearly seen that the feature map tends to characterize the global information after adding the attention mechanism, but still does not extract the hand features well for the occluded regions and hand interaction edges. The SSA module, on the other hand, aggregates more hand information and performs better at the interaction edges, but it will characterize the occluded part as the hand, which introduces interference information to a certain extent. The SSCA module expands the range of the receptive field through the spatial plane selection mechanism, and then directionally strengthens the hand region features through the selection of the channel planes. Furthermore, by benefiting from linear directional scanning, the occluded areas are characterized using adjacent hand information, which reduces the introduction of interference through this global mapping capability.
In Table 6, it can be found that after the introduction of the state selection of the channel plane, the number of parameters of the model is only slightly increased by 0.01M, however, a significant improvement of nearly 2% is realized in the two performance metrics of PA-MPJPE and PA-MPVPE. This is due to the fact that the state selection in the channel plane enhances the selection and fusion of image features, and is able to extract more important channel information. The current state-space model focuses more on multi-directional scanning in the spatial plane, which is the reason for the 0.3% improvement in the F@5 metrics when using SSA. Using only the spatial plane for state selection, each pixel will focus more on neighboring pixels. However, observing other metrics will show that the channel dimension in the image will then contain richer feature information, which happens to be ignored by SSA.
Figure 4: Qualitative results for the three datasets. From left to right are the input image containing severe occlusions and angular distortions, the attention mapping map obtained by SSCA, the heat map of keypoints in Regressor, and the hand mesh.
Figure 5: Attention characterization map. It is worth noting that the SSCA module is able to obtain sufficient hand information in the hand interaction edge region and the occlusion region. And the method expands the range of the receptive field so that it can capture the global feature mapping. (Orange is the occluded region and red is the hand interaction edge region)
5 Conclusion
In order to effectively mitigate the challenges of the hand pose estimation task under occlusion conditions, we introduce state-space modeling into the field of hand pose estimation for the first time and propose the HandSSCA method. The method expands the effective receptive field and acquires hand features under occlusion by parallel cross-scanning in the spatial plane and the channel plane. Among them, the spatial plane is able to acquire feature associations in different directions for each pixel, and the channel plane activates the target region features through state selection. This novel attention architecture is able to acquire sufficient global information and expand the effective sensory field while keeping the number of parameters small. Extensive experimental results show that this method is able to accurately predict hand postures from heavily occluded situations. Meanwhile, we expect this spatial and channel-parallel state-selective attention architecture to make progress in different downstream tasks.